Chapter 15, Problem 036 GO Under constant-volume conditions, 3000 J of heat is added to 1.3 moles of an ideal gas. As a resulhelp please
Answer #1

For O 1 constant volume: Qu= ncr DT Op = nx Cpx AT = nx (Cr+R) AT = ncrat & NRAT - Qv+ NROT = 3000 + 1-3x 8-314x185 = 4999.5